Gender Selection Assume that the groups consists of 16 couples.

a. Find the mean and standard deviation for the numbers of girls in groups of 16 births.
Answer #1

Ans:

Probability of being a girl,p=0.5

n=16

mean=np=16*0.5=8

standard deviation=sqrt(16*0.5*(1-0.5))=2